# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

### v1

@@ Appendix A Spectrum of the central source
As one can see in Fig. 6, both blue-shifted and red-shifted lines of highly ionized (H-like and He-like) heavy elements (in particular, silicon, sulfur, iron, and nickel) moving at 0.26 speed of light can be easily discriminated and identified. They arise from the approaching (with apparent redshift -0.085) and the receding jet (with apparent redshift 0.167), respectively. In agreement with the findings of numerous previous observations (see Medvedev et al. 2019, for a recent summary), a factor of 10 overabundance of nickel (with respect to iron) in the gas of the jets is required to describe the spectrum well (most clearly indicated by the ratio of the Ni XXVII and Fe XXV lines). Spectral resolution of eROSITA is not sufficient to resolve the width of the lines, which is known to be dominated by the ballistic motion of gas within a cone of with opening angle , as measured by high resolution spectroscopy with Chandra gratings (e.g. Marshall et al. 2002). Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev & Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al. 2018, 2019). Comparison of the spectra taken in the two epochs of observations separated by one orbital period shows very good consistency between them333Since one orbital period separation is small compared to the precession period and the observations were close to an extreme point of the precession curve, the change in the Doppler shifts of the jets is small., with the biggest difference observed in the 6-7 keV band. This might be related to the change in the visible geometry of this region or to the physical conditions within it (see a discussion in Medvedev et al. 2019). High-resolution spectroscopy of this emission with the XRISM observatory (Tashiro et al. 2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin & Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al. 1988; Brinkmann & Kawai 2000; Khabibullin et al. 2016). The SS433’s spectrum presented here illustrates the spectroscopic capabilities of eROSITA and shows that, for sufficiently bright sources, reliable continuum and line characterization can be achieved up to 10 keV.

### v2

@@ Appendix A Spectrum of the central source
As one can see in Fig. 6, both blue-shifted and red-shifted lines of highly ionized (H-like and He-like) heavy elements (in particular, silicon, sulfur, iron, and nickel) moving at 0.26 speed of light can be easily discriminated and identified. They arise from the approaching (with apparent redshift -0.085) and the receding jet (with apparent redshift 0.167), respectively. In agreement with the findings of numerous previous observations (see Medvedev et al. 2019, for a recent summary), a factor of 10 overabundance of nickel (with respect to iron) in the gas of the jets is required to describe the spectrum well (most clearly indicated by the ratio of the Ni XXVII and Fe XXV lines). Spectral resolution of eROSITA is not sufficient to resolve the width of the lines, which is known to be dominated by the ballistic motion of gas within a cone of with opening angle , as measured by high resolution spectroscopy with Chandra gratings (e.g. Marshall et al. 2002). Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev and Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al. 2018, 2019). Comparison of the spectra taken in the two epochs of observations separated by one orbital period shows very good consistency between them333Since one orbital period separation is small compared to the precession period and the observations were close to an extreme point of the precession curve, the change in the Doppler shifts of the jets is small., with the biggest difference observed in the 6-7 keV band. This might be related to the change in the visible geometry of this region or to the physical conditions within it (see a discussion in Medvedev et al. 2019). High-resolution spectroscopy of this emission with the XRISM observatory (Tashiro et al. 2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin and Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al. 1988; Brinkmann and Kawai 2000; Khabibullin et al. 2016). The SS433’s spectrum presented here illustrates the spectroscopic capabilities of eROSITA and shows that, for sufficiently bright sources, reliable continuum and line characterization can be achieved up to 10 keV.